OpenStudy (anonymous):

Find the exact solutions of x^2 - (y - 12)^2 = 144 and y = -x^2 and.. solve the system: y^2 -4x^2 and y = 2x

OpenStudy (anonymous):

for the first one, solve \[x^2-(-x^2-12)^2=144\]

OpenStudy (anonymous):

which is the same as \[x^2-(x^2+12)^2=144\] this will have no real zeros, so my guess is there might be a typo in the question

OpenStudy (anonymous):

could it be \(x^2 +(y - 12)^2 = 144\) ?

OpenStudy (anonymous):

I got the answer to the second question - there's not a typo I don't think I'm just not sure what it's asking for?

OpenStudy (anonymous):

In the first question that is

OpenStudy (anonymous):

there is no real answer for x^2 - (y - 12)^2 = 144 and y = -x^2

OpenStudy (anonymous):

so what would i write then? no solution? i know i tried to solve the equation myself and it turned out funky.

OpenStudy (anonymous):

no real solution

OpenStudy (anonymous):

for the second one I got no solution as well, do you agree?

OpenStudy (anonymous):

well 0=4, but basically no solution.